8=56
7=42
6=30
5=20
3=?

Found 2 solutions by checkley79, acubed1234:
Answer by checkley79(3341)

8=56 8*7=56
7=42 7*6=42
6=30 6*5=30
5=20 5*4=20
3=? 3*2=6 ans.

Answer by acubed1234(5)


I wrote out a general solution here . . . https://www.scribd.com/doc/260182194/Elementary-Sequences